# TI-DeepONet: Learnable Time Integration for Stable Long-Term Extrapolation

### v1

@@ C.1 One-dimensional Burgers’ Equation
As previously discussed, the learnable slope coefficients $\boldsymbol{\alpha}=[\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}]$ are predicted by the additional MLP, which takes the current solution state $u^{i}$ as input, i.e., $\boldsymbol{\alpha}=\text{MLP}_{\theta}(u^{i})$. Analyzing the evolution of these solution-specific coefficients over training epochs provides insight into how the network adapts the time integrator to local dynamics, in contrast to using a fixed set of RK4 coefficients. This comparison is illustrated in Fig. A1. The dashed blue line in each plot indicates the reference slope coefficients typically used in the RK4 method, derived from a Taylor series expansion. These coefficients result in a fourth-order accurate scheme and are generally expected to provide high accuracy for traditional time integrators. However, in the context of the operator learning framework, these coefficients also adapt to the approximation error of the neural operator. Making these slope coefficients learnable allows the model to better compensate for such errors, thereby improving the overall accuracy of the framework.
@@ C.2 One-dimensional Korteweg-de Vries (KdV) Equation
where $sech$ is hyperbolic secant, $P$ is the period in space, % is the modulo operator, $i=\{1,2\}$, $k\in[0.5,1]$ and $d\in[0,1]$ are the coefficients that determine the height and location of the soliton peak, respectively. The dataset is curated by generating $N=1000$ distinct initial conditions and their temporal evolution is modeled using the midpoint method. The 1D spatial domain, $\Omega=[0,10]$, is discretized into 100 uniformly spaced grid points, resulting in a spatial resolution of $\Delta x=0.1$. For the temporal domain, we consider a uniform grid of 201 points spanning the time domain $t\in[0,5]$, which yields a $\Delta t=0.025$.
@@
To assess the extrapolation accuracy of the different methods, we train only on the first half of the temporal domain, i.e., $t_{\text{train}}\in[0,2.5]$, and evaluate the predictive performance over the entire domain ($t\in[0,5]$). To reiterate, in the full rollout setting, the inputs to the trunk network are constructed from a meshgrid of $(x,t_{\text{train}})$, whereas in the autoregressive setting, only the spatial coordinates $(x)$ are used as trunk inputs. Moreover, while the initial condition samples serve as the input to the branch network in the DeepONet full rollout model, the input to the branch network in the autoregressive and TI-based DeepONet frameworks consists of the stacked solution states from $t=0$ to $t=2.5$, i.e., $[u^{0}(x),u^{1}(x),\dots,u^{100}(x)]$. The train-test split is kept at 0.8.
@@
Table A4 presents the network architecture details, including the number of layers, neurons per layer, activation functions, total epochs to convergence, and batch size. Similarly, for the TI(L)-DeepONet framework, we employ a secondary feedforward network to predict the four RK4 slope coefficients. This network consists of two hidden layers with 64 and 48 neurons, respectively, followed by an output layer with four neurons. The hidden layers use the tanh activation function, while a softmax activation is applied at the output layer to ensure that the predicted coefficients sum to one. Training was performed using the Adam optimizer with an initial learning rate of $10^{-3}$, which was exponentially decayed by a factor of 0.95 every 5000 epochs. A similar learning rate schedule was also used for training the auxiliary network tasked with learning the RK4 slope coefficients.
Table A4: 1D KdV equation: Neural network architecture and training settings across different methods.
@@
TI(L)-DeepONet [128]$\times$6, [80] [128]$\times$6, [80] Tanh $1.5\times 10^{5}$ 256
TI-DeepONet [128]$\times$6, [80] [128]$\times$6, [80] Tanh $1\times 10^{5}$ 256
Full rollout [150, 250, 450, 380, 320, 300, 80] [200, 220, 240, 250, 260, 280, 300, 80] Swish $1\times 10^{5}$ 256
Autoregressive [128]$\times$6, [80] [128]$\times$6, [80] Tanh $1\times 10^{5}$ 256
C.3 Two-dimensional Burgers’ Equation
Data generation
In the 2D Burgers’ transport dynamics problem, we consider periodic boundary conditions, where the solution $u(t,x,y)$ is a scalar field, and the initial condition $u_{0}(x,y)$ is sampled from a periodic Matérn-type GRF with length scale $l=0.125$ and standard deviation $\sigma=0.15$.
Data preparation
A total of $N=1000$ initial condition functions were generated. Each initial condition was discretized on a $32\times 32$ spatial grid, and the corresponding solution field was evolved over $101$ uniformly spaced time steps, resulting in a spatiotemporal output field of size $101\times 32\times 32$ for each of the 1000 initial condition realizations. The training and testing dataset were prepared using the same strategy as in the previous two cases.
In the full rollout setting, our aim is to learn a solution operator, $\mathcal{G}_{\boldsymbol{\theta}}$, that performs the mapping: $u_{0}(x,y)\rightarrow u(t,x,y)$. Using a train-test split of 0.8, 800 realizations of initial conditions and their corresponding ground truth solutions were used for training, with 200 samples of initial conditions and ground truth solution fields kept aside for testing. The trunk network takes as input a triplet of spatiotemporal coordinates, $(t,x,y)$; in a similar spirit to the previous cases, we only provide the first 33 out of 101 time steps during training. Thus, the input meshgrid for the trunk network is constructed using one-third of the temporal domain, i.e., $(t,x,y),$ where $t=t_{\text{train}}\in[0,0.33]$. For the autoregressive and TI-based frameworks, the input to the branch network consists of stacked solution states from $t=0$ to $t=0.33$, i.e., the dataset $[u^{0}(x,y),u^{1}(x,y),\dots,u^{33}(x,y)]$. The trunk network in this case takes only the spatial coordinates $(x,y)$ as input.
@@ C.3 Two-dimensional Burgers’ Equation
Table A5 presents the network architecture details, including the number of layers and neurons per layer, activation functions, total epochs to convergence, and batch size. In this case, for the TI(L)-DeepONet architecture, we employ a similar combination of 2D convolutional blocks and MLP layers as used in the branch network to learn the RK4 slope coefficients. The convolutional blocks utilize the same combination of Conv2D and pooling layers as the branch, with the only difference being that the Conv2D layers have a reduced number of filters (32 instead of 64). This is followed by a flattening layer, after which the intermediate representation is passed through two fully connected hidden layers, each containing 32 neurons, and finally to an output layer with 4 neurons that predicts the RK4 slope coefficients.
The Conv2D layers use a ReLU activation function, the fully connected hidden layers use tanh, and the output layer applies a softmax activation to ensure that the predicted slope coefficients sum to one. Training was performed using the Adam optimizer with an initial learning rate of $10^{-3}$, which was exponentially decayed by a factor of 0.95 every 5000 epochs. The auxiliary network responsible for predicting the learnable RK4 slope coefficients was trained with a similar exponential decay schedule, but with a higher initial learning rate of $5\times 10^{-3}$.
Table A5: 2D Burgers’ equation: Neural network architecture and training settings across different methods.
@@
TI(L)-DeepONet Conv2D(features = 64, kernel size = (3,3), strides = 1), MaxPool(window shape=(2, 2), strides = (2, 2)), Conv2D(features = 64, kernel size = (2,2), strides = 1), AvgPool(window shape = (2,2), strides = (2,2)), MLP([256, 128, 100]) [128]$\times$4, [100] Conv2D: ReLU, MLP: Tanh $1.5\times 10^{5}$ 64
TI-DeepONet Conv2D(features = 64, kernel size = (3,3), strides = 1), MaxPool(window shape=(2, 2), strides = (2, 2)), Conv2D(features = 64, kernel size = (2,2), strides = 1), AvgPool(window shape = (2,2), strides = (2,2)), MLP([256, 128, 100]) [128]$\times$4, [100] Conv2D: ReLU, MLP: Tanh $1.2\times 10^{5}$ 64
Full rollout Conv2D(features = 64, kernel size = (3,3), strides = 1), MaxPool(window shape=(2, 2), strides = (2, 2)), Conv2D(features = 64, kernel size = (2,2), strides = 1), AvgPool(window shape = (2,2), strides = (2,2)), MLP([256, 128, 100]) [128]$\times$3, [100] Conv2D: ReLU, MLP: SiLU $1\times 10^{5}$ 128
Autoregressive Conv2D(features = 64, kernel size = (3,3), strides = 1), MaxPool(window shape=(2, 2), strides = (2, 2)), Conv2D(features = 64, kernel size = (2,2), strides = 1), AvgPool(window shape = (2,2), strides = (2,2)), MLP([256, 128, 100]) [128]$\times$6, [100] Conv2D: ReLU, MLP: SiLU $1\times 10^{5}$ 64
Figure A2: 2D Burgers’ equation: Solution field for a representative sample, shown across both the training regime ($t\in[0,0.33]$) and the extrapolation regime ($t\in[0.33,1]$), for all frameworks. Corresponding error plots are presented in Figure 6.

### v2

@@ C.1 One-dimensional Burgers’ Equation
As previously discussed, the learnable slope coefficients $\boldsymbol{\alpha}=[\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}]$ are predicted by the additional NN, which takes the current solution state $u^{i}$ as input, i.e., $\boldsymbol{\alpha}=\text{NN}_{\theta}(u^{i})$. Analyzing the evolution of these solution-specific coefficients over training epochs provides insight into how the network adapts the time integrator to local dynamics, in contrast to using a fixed set of RK4 coefficients. This comparison is illustrated in Fig. A1. The dashed blue line in each plot indicates the reference slope coefficients typically used in the RK4 method, derived from a Taylor series expansion. These coefficients result in a fourth-order accurate scheme and are generally expected to provide high accuracy for traditional time integrators. However, in the context of the operator learning framework, these coefficients also adapt to the approximation error of the neural operator. Making these slope coefficients learnable allows the model to better compensate for such errors, thereby improving the overall accuracy of the framework.
@@ C.2 One-dimensional Korteweg-de Vries (KdV) Equation
where sechsech is hyperbolic secant, $P$ is the period in space, % is the modulo operator, $i=\{1,2\}$, $k\in[0.5,1]$ and $d\in[0,1]$ are the coefficients that determine the height and location of the soliton peak, respectively. The dataset is curated by generating $N=1000$ distinct initial conditions and their temporal evolution is modeled using the midpoint method. The 1D spatial domain, $\Omega=[0,10]$, is discretized into 100 uniformly spaced grid points, resulting in a spatial resolution of $\Delta x=0.1$. For the temporal domain, we consider a uniform grid of 201 points spanning the time domain $t\in[0,5]$, which yields a $\Delta t=0.025$.
@@
To assess the extrapolation accuracy of the different methods, we train only on the first half of the temporal domain, i.e., $t_{\text{train}}\in[0,2.5]$, and evaluate predictive performance over the entire domain ($t\in[0,5]$). To reiterate, in the full rollout setting, the inputs to the trunk network are constructed from a meshgrid of $(x,t_{\text{train}})$, whereas in the autoregressive setting, only the spatial coordinates $(x)$ are used as trunk inputs. Similarly, the same meshgrid that serves as trunk inputs during DeepONet FR training becomes the axes along which the 2D Fourier transform is performed (all $x$ coordinates, seen $t$ coordinates) for FNO FR training. While the initial condition samples serve as input to the branch network in the DeepONet full rollout model, the input to the branch network in the autoregressive and TI-based DeepONet frameworks consists of stacked solution states from $t=0$ to $t=2.5$, i.e., $u^{0}(x),u^{1}(x),\ldots,u^{100}(x)]$. For autoregressive FNO, the inputs and outputs are the solution pairs $u^{i}(x)$ and $u^{i+1}(x)$, with the 1D Fourier transform applied along the spatial dimension $x$ only. The train-test split is maintained at 0.8.
@@
Tables A5 and A6 present the network architecture details, including the number of layers, neurons per layer, activation functions, total epochs to convergence, batch size, and Fourier modes (where applicable). For the TI(L)-DeepONet framework, we employ a secondary feedforward network to predict the four RK4 slope coefficients. This network consists of two hidden layers with 64 and 48 neurons, respectively, followed by an output layer with four neurons. The hidden layers use the tanh activation function, while a softmax activation is applied at the output layer to ensure that the predicted coefficients sum to one. Training was performed using the Adam optimizer with an initial learning rate of $10^{-3}$, which was exponentially decayed by a factor of 0.95 every 5000 epochs. A similar learning rate schedule was used for training the auxiliary network tasked with learning the RK4 slope coefficients. The same learning rate schedule with identical hyperparameters was employed for FNO FR training. For FNO AR, an exponential decay learning rate schedule was also used, but with an initial learning rate of $10^{-3}$ that decayed by a factor of 0.96 every 2000 steps.
Table A5: 1D KdV equation: Neural network architecture and training settings across different DeepONet methods.
@@
[…]
DON FR [150, 250, 450, 380, 320, 300, 80] [200, 220, 240, 250, 260, 280, 300, 80] Swish $1\times 10^{5}$ 256
[…]
Table A6: 1D KdV equation: Neural network architecture and training settings across different FNO-based methods.
Method Fourier Blocks Modes Hidden Dim Activation Epochs Batch Size
FNO AR 6 24 64 GELU 30000 128
FNO FR 6 32 64 GELU 10000 64
C.3 One-dimensional Kuramuto-Sivashinsky (KS) Equation
Data generation:
We consider the 1D Kuramoto-Sivashinsky (KS) equation on a spatial domain $x\in[0,L]$ with $L=6\pi$, equipped with periodic boundary conditions. The initial conditions $u(x,0)=s(x)$ are sampled from mean-zero Gaussian random fields (GRFs) with periodic boundary conditions. Specifically, we generate GRFs with covariance operator $\mathcal{C}=\sigma^{2}(-\Delta+\tau^{2}I)^{-\gamma}$, where $\gamma=1$, $\tau=2$, and $\sigma=1$. The spatial domain is discretized using $N_{x}=128$ grid points, and the temporal domain $t\in[0,30]$ is discretized with $N_{t}=300$ timesteps. The PDE is solved using the exponential time-differencing fourth-order Runge-Kutta (ETDRK4) method with a timestep of $\Delta t=0.1$. We generate a total of 3000 samples, each consisting of the initial condition and the corresponding spatiotemporal solution field $u(x,t)$. The dataset is partitioned using a train-test split ratio of 0.8, yielding $N_{\text{train}}=2400$ training samples and $N_{\text{test}}=600$ testing samples. The periodic boundary conditions ensure that the solution remains in the space of mean-zero periodic functions throughout the evolution, preserving the chaotic dynamics characteristic of the KS equation.
Data preparation:
For evaluating the extrapolation capabilities of various methods, we restrict training to the initial half of the temporal domain, specifically $t_{\text{train}}\in[0,15]$, while testing spans the complete temporal interval $t\in[0,30]$. In the full rollout configuration, the trunk network receives spatiotemporal coordinate pairs $(x,t_{\text{train}})$ arranged in a meshgrid format. Conversely, the autoregressive approach utilizes only spatial coordinates $x$ as trunk inputs. For FNO FR, the 2D Fourier transform operates along both spatial and temporal dimensions using the same coordinate grid employed in DeepONet FR training. The branch network architecture differs between approaches: full rollout DeepONet processes initial conditions $s(x)$ directly, while autoregressive and TI-based variants receive a sequence of solution snapshots spanning $t=0$ to $t=15$, comprising states $u^{0}(x),u^{1}(x),\ldots,u^{150}(x)]$, with the corresponding output solution states $u^{1}(x),u^{2}(x),\ldots,u^{151}(x)]$. The autoregressive FNO framework also operates on consecutive solution pairs $(u^{i}(x),u^{i+1}(x))$ as input-output mappings, employing a 1D Fourier transform exclusively in the spatial dimension.
Training architecture:
Tables A7 and A8 present the network architecture details, including the number of layers, neurons per layer, activation functions, total epochs to convergence, batch size, and Fourier modes (where applicable). All DeepONet variants employ Fourier feature encoding in the trunk network. For the full rollout DeepONet, both spatial and temporal coordinates $(x,t)$ are augmented with sinusoidal features computed as $\sin(\pi kx),\cos(\pi kx),\sin(\pi kt),\cos(\pi kt)]$ for $k\in\{1,2,\ldots,10\}$, expanding the input dimension from 2 to 42. For the autoregressive and TI-based variants, only the spatial coordinate $x$ is augmented with Fourier features $\sin(\pi kx),\cos(\pi kx)]$, expanding the input dimension from 1 to 21. For the TI(L)-DeepONet framework, we employ an auxiliary feedforward network to predict the four RK4 slope coefficients. This network consists of two hidden layers with 64 neurons each, followed by an output layer with four neurons. The hidden layers utilize the tanh activation function, while a softmax activation is applied at the output layer to ensure the predicted coefficients sum to one. All DeepONet variants (DON FR, DON AR, TI-DON, and TI(L)-DON) are trained using the AdamW optimizer with an initial learning rate of $10^{-3}$, exponentially decayed by a factor of 0.95 every 5000 epochs, and a weight decay of $10^{-4}$. For TI(L)-DeepONet specifically, the auxiliary RK4 network is trained separately using the Adam optimizer with an initial learning rate of $2\times 10^{-3}$ and the same exponential decay schedule. The FNO variants employ the Adam optimizer with an initial learning rate of $10^{-3}$, exponentially decayed by a factor of 0.96 every 2000 steps for both FNO FR and FNO AR configurations.
Table A7: 1D KS equation: Neural network architecture and training settings across different DeepONet methods.
Method Branch Net Trunk Net Activation Epochs Batch Size
[…]
Table A8: 1D KS equation: Neural network architecture and training settings across different FNO-based methods.
Method Fourier Blocks Modes Hidden Dim Activation Epochs Batch Size
FNO AR 6 64 64 GELU 10000 64
FNO FR 6 64 64 GELU 7000 64
C.4 Two-dimensional Inviscid Burgers’ Equation
Data generation:
For the 2D inviscid Burgers’ equation, we consider periodic boundary conditions on a unit square domain $(x,y)\in[0,1]^{2}$. The initial conditions $u(x,y,0)=s(x,y)$ are sampled from periodic Matérn-type Gaussian random fields with length scale $l=0.1$ and standard deviation $\sigma=0.2$. The spatial domain is discretized using a $32\times 32$ grid with $N_{x}=N_{y}=32$ points. The PDE is solved using a fourth-order Runge-Kutta scheme with timestep $\Delta t=10^{-4}$ up to final time $t=1.0$. Solutions are saved at intervals of $\Delta t_{\text{save}}=0.01$, resulting in 101 temporal snapshots. The kinematic viscosity is set to $\nu=10^{-4}$, which is sufficiently small to approximate inviscid behavior while maintaining numerical stability. We generate a total of 1000 samples, each consisting of the initial condition and the corresponding spatiotemporal solution field $u(x,y,t)$ of size $101\times 64\times 64$.
Data preparation:
The dataset is partitioned using a train-test split ratio of 0.8, yielding $N_{\text{train}}=800$ training samples and $N_{\text{test}}=200$ testing samples. To evaluate extrapolation capabilities, we train only on the first third of the temporal domain, i.e., $t_{\text{train}}\in[0,0.33]$ (33 out of 101 time steps), and assess performance over the entire domain $t\in[0,1.0]$. This represents the most challenging extrapolation scenario in our study, requiring $3\times$ temporal extrapolation with no exposure to shock formation during training.
For full rollout methods, our aim is to learn a solution operator $\mathcal{G}_{\boldsymbol{\theta}}$ that maps initial conditions to the full solution: $u_{0}(x,y)\rightarrow u(t,x,y)$. The trunk network receives $(x,y,t_{\text{train}})$ coordinates as input, while the branch network processes the initial conditions. For FNO FR, the 3D Fourier transform is performed along all three dimensions $(x,y,t_{\text{train}})$, capturing the full spatiotemporal frequency content within the training domain. In contrast, autoregressive and TI-based methods use only spatial coordinates $(x,y)$ for the trunk network input. Their branch networks receive stacked solution states $u^{0}(x,y),u^{1}(x,y),\ldots,u^{33}(x,y)]$ from the training temporal window. FNO AR employs a 2D Fourier transform along the spatial dimensions $(x,y)$ only, consistent with its autoregressive nature to learn a one-step mapping between pairs of consecutive solution states $(u^{i}(x,y),u^{i+1}(x,y))$. Note that for the autoregressive/TI-based variants, explicit temporal information is absent and instead is assimilated in the batch dimension.
@@ C.4 Two-dimensional Inviscid Burgers’ Equation
Tables LABEL:tab:2d_burgers_don-architecture and A10 presents the network architecture details, including the number of layers and neurons per layer, activation functions, total epochs to convergence, batch size and Fourier modes (if applicable). All DeepONet variants employ Fourier feature encoding in the trunk network. For the full rollout DeepONet, the spatial and temporal coordinates $(x,y,t)$ are augmented with sinusoidal features computed as $\sin(\pi kx),\cos(\pi kx),\sin(\pi ky),\cos(\pi ky),\sin(\pi kt),\cos(\pi kt)]$ for $k\in\{1,2,\ldots,10\}$, expanding the input dimension from 3 to 63. For the autoregressive and TI-based variants, only the spatial coordinates $(x,y)$ are augmented with Fourier features $\sin(\pi kx),\cos(\pi kx),\sin(\pi ky),\cos(\pi ky)]$, expanding the input dimension from 2 to 42. In this case, for the TI(L)-DeepONet architecture, we employ a similar combination of 2D convolutional blocks and MLP layers as used in the branch network to learn the RK4 slope coefficients. The convolutional blocks utilize the same combination of Conv2D and pooling layers as the branch, with the only difference being that the Conv2D layers have a reduced number of filters (32 instead of 64). This is followed by a flattening layer, after which the intermediate representation is passed through two fully connected hidden layers and finally to an output layer with 4 neurons that predicts the RK4 slope coefficients.
The Conv2D layers use a ReLU activation function, the fully connected hidden layers use tanh, and the output layer applies a softmax activation to ensure that the predicted slope coefficients sum to one. Training was performed using the Adam optimizer with an initial learning rate of $10^{-3}$, which was exponentially decayed by a factor of 0.95 every 5000 steps. The auxiliary network responsible for predicting the learnable RK4 slope coefficients was trained with a similar exponential decay schedule, but with a higher initial learning rate of $2\times 10^{-3}$. For both FNO AR and FNO FR variants, we employ a similar exponential decay schedule as the DeepONet variants albeit with a decay factor of 0.96 every 2000 steps.
Table A9: 2D Inviscid Burgers’ equation: Neural network architecture and training settings across different methods.
@@
[…]
Table A10: 2D Burgers’ equation: Neural network architecture and training settings across different FNO-based methods.
Method Fourier Blocks Modes Hidden Dim Activation Epochs Batch Size
FNO AR 6 12 64 GELU 1000 128
FNO FR 6 12 64 GELU 5000 32
Figure A2: 2D Inviscid Burgers’ equation: Solution field for a representative sample, shown across both the training regime ($t\in[0,0.33]$) and the extrapolation regime ($t\in[0.33,1]$), for all frameworks. Corresponding error plots are presented in Figure 7.